Two pharma companies, namely Zydus and Torrent pollute the Sabarmati River. They each dump

100 tons of carbon mixed water into the river. The cost of reducing production of carbon mixed

water for Zydus is Rs. 100 per ton, whereas Torrent will have to spend Rs. 1000 per ton of carbon

mixed water to reduce pollution. The Ahmedabad Municipal Corporation wants to reduce carbon

mixed water generation to 50 tons, from the current 200 tons.

(i) If AMC knew their costs of reduction for each company, what reductions would it impose to reach

its overall goal? What would be the cost to each company, and what would be the total cost to the

companies together?



1
Expert's answer
2021-11-12T08:00:09-0500

The Ahmedabad Municipal Corporation would impose a reduction of 100 tons for  Zydus and the

remaining 50 tons for Torrent . This solution would minimize the total

cost of reducing the pollution. This cost would be:


"100\times 100 + 1000 \times 50\\
 = 10000 + 50000 = 60000"


Rs. 60000